# Understanding and Improving Shampoo and SOAP via Kullback–Leibler Minimization

### v3

@@ 1 Introduction
In this work, we introduce a KL perspective that interprets the estimation schemes of Shampoo and SOAP as solutions to KL-minimization problems for covariance estimation. Our approach naturally extends to tensor-valued settings, where some existing theoretical interpretations may not apply. This perspective reveals a key limitation (illustrated in Fig.˜5): the Kronecker-structured estimators used by Shampoo and SOAP do not adequately solve the corresponding KL-minimization problem. This limitation, in turn, opens new opportunities for improvement. Leveraging this insight, we refine the estimation rules of Shampoo and SOAP and develop practical KL-based schemes—KL-Shampoo and KL-SOAP—that meet or exceed the performance of Shampoo and SOAP for NN (pre-)training while maintaining SOAP-level per-iteration runtime. Notably, KL-Shampoo does not rely on Adam to achieve competitive performance, thereby avoiding Adam’s additional memory overhead (Footnote˜1). In addition, our practical techniques developed for KL-Shampoo (Sec.˜4) can be adapted to strengthen existing Shampoo variants and make the trace-scaling variant competitive without step-size grafting (Fig.˜10, Appx.˜H), while also achieving SOAP-level per-iteration runtime (Fig.˜10, Appx.˜H). Empirically, we show that KL-based methods are competitive for training a range of NNs and remain as flexible as Shampoo and SOAP for tensor-valued weights. Surprisingly, KL-Shampoo consistently outperforms the other methods in our experiments (Figs.˜1 and 5). Overall, our KL-based approach provides a principal way for designing structured methods in NN optimization.
@@ 2 Background
where $\hat{{\bm{g}}}:={\bm{Q}}^{\top}{\bm{g}}=\operatorname{vec}({\bm{Q}}_{a}^{\top}{\bm{G}}{\bm{Q}}_{b})$ is a “projected" gradient vector in eigenbasis $\bm{Q}$} and recall that ${\bm{S}}:={\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$ is SOAP’s preconditioner. Here, we leverage the Kronecker structure and orthogonality of the eigenbasis to move from the left to the right in the last line of Eq.˜2. Note that this EMA weight $\beta_{2}$ is defined as $1-\beta_{2}^{\text{(Adam)}}$, where $\beta_{2}^{\text{(Adam)}}$ is Adam’s (RMSProp’s) $\beta_{2}$. We use this definition rather than Adam’s because we want to further interpret this moving-average scheme through the lens of our KL perspective.
@@ 3 Second Moment Estimation via Kullback–Leibler Minimization
where $\smash{\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]}$ and $\bm{S}$} are considered as Gaussian’s covariance, $\mathrm{det}(\cdot)$ denotes the matrix determinant of its input, and $\kappa\!\geq\!0$ is a damping weight to ensure the positive-definiteness of $\smash{\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\!+\!\kappa{\bm{I}}}$ if necessary. Mathematically, this KL divergence coincides (up to a factor of $\nicefrac{{1}}{{2}}$) with the log-determinant divergence widely used in matrix optimization (Dhillon & Tropp, 2008; Kulis et al., 2009; Sra, 2016), which is defined for any pair of SPD matrices and does not require a zero-mean assumption. This additional zero-mean Gaussian viewpoint provides a probabilistic interpretation of this SPD-aware “distance”, even when the target matrix is not itself a second moment, such as the curvature matrix used in quasi-Newton methods (Fletcher, 1991; Waldrip & Niven, 2016). Moreover, the KL divergence naturally extends to tensor-valued cases, such as a 3D tensor gradient, ${\bm{G}}\in\mathcal{R}^{d_{a}\times d_{b}\times d_{c}}$, by considering a structured preconditioner ${\bm{S}}={\bm{S}}_{a}\otimes{\bm{S}}_{b}\otimes{\bm{S}}_{c}$ to approximate the flattened gradient second moment, $\smash{\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]}$, where matrix ${\bm{S}}_{k}\in\mathcal{R}^{d_{k}\times d_{k}}$ is SPD for $k\in\{a,b,c\}$.
@@
Many works (Morwani et al., 2025; Eschenhagen et al., 2025; An et al., 2025; Xie et al., 2025) primarily focus on matrix-valued weights and interpret Shampoo’s and SOAP’s estimation rules from the Frobenius-norm perspective. However, this norm does not account for the SPD constraint implicitly imposed on Shampoo’s and SOAP’s preconditioners, which ensures that the preconditioned gradient direction is a descent direction (Nesterov et al., 2018). As emphasized in the literature (Pennec et al., 2006; Bhatia, 2007), it is more appropriate to consider a “distance” that respects this constraint. We adopt the KL divergence because it naturally incorporates the SPD constraint, is widely used for covariance estimation (Amari, 2016; Minh & Murino, 2017), and provides a unified framework for reinterpreting and improving Shampoo’s estimation—even in tensor-valued settings where some existing interpretations based on singular value decomposition (Van Loan & Pitsianis, 1993) may not apply. In numerical optimization, the KL divergence, known as a merit function (Byrd & Nocedal, 1989), offers a unifying interpretation (Fletcher, 1991; Waldrip & Niven, 2016) and extension (Kanamori & Ohara, 2013a, b) of the low-rank estimation schemes of BFGS and DFP. In contrast, the standard Frobenius norm cannot recover these updates without additional weighting (see Sec. 6.1 of Nocedal & Wright (2006)). In statistical estimation and inference, this KL divergence is also preferred over the Frobenius norm (James et al., 1961; Kivinen & Warmuth, 1999; Davis & Dhillon, 2006; Khan & Lin, 2017; Lin et al., 2019; Kunstner et al., 2021).
@@ 3.2 Improving Shampoo’s estimation: Idealized KL-Shampoo
˜2 (proof in Appx.˜B) establishes a closed-form condition (see Eq.˜4) when simultaneously learning both Kronecker factors to minimize the KL problem. Statistically, this condition corresponds to the maximum-likelihood covariance estimator of a zero-mean matrix Gaussian (Dutilleul, 1999). This condition implies that $\bm{G}$} is generated from a zero-mean matrix Gaussian with row-wise covariance ${\bm{S}}_{a}^{*}$ and column-wise covariance ${\bm{S}}_{b}^{*}$. Under this condition, Shampoo-style preconditioning naturally induces matrix-Gaussian (row and column) whitening. Our KL-based approach further extends this to tensor-Gaussian whitening for tensor-valued gradients—without the prohibitive cost typically associated with SVD-based methods. Notably, this kind of whitening arises from minimizing KL divergence, not the Frobenius norm. In machine learning, Lin et al. (2019, 2024) treated this condition as a theoretical example of a multilinear exponential-family (see Sec 5 of Lin et al. (2019)) for Kronecker-based optimization via natural gradient descent on an exponential-family manifold, while more recently, Vyas et al. (2025b) considered a similar condition motivated heuristically by gradient whitening. However, we cannot directly use this condition due to the correlated update between ${\bm{S}}_{a}$ and ${\bm{S}}_{b}$. For example, solving ${\bm{S}}_{a}^{*}$ requires knowing ${\bm{S}}_{b}^{*}$ in Eq.˜4 or vice versa. In practice, this optimal condition is often unachievable because the expectations in Eq.˜4 have to be approximated. Thus, we consider an estimated ${\bm{S}}_{k}$ to approximate ${\bm{S}}_{k}^{*}$ for $k\in\{a,b\}$ and propose an exponential moving average (EMA) scheme in practice:
@@ 3.3 Comparison with Kronecker Schemes Using Alternative Divergences
Morwani et al. (2025) consider a two-sided Shampoo variant based on the Frobenius norm and derive the optimal solution via rank-1 singular value decomposition (SVD) of the outer product $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ (Van Loan & Pitsianis, 1993). However, this optimal solution is often unattainable in practice and is computationally expensive for two reasons: (1) the expectation $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ must be approximated; and (2) performing the SVD itself is costly—yielding complexity ($O(d_{a}^{2}d_{b}^{2})$) in general even for rank-1 SVD—which is higher than the eigendecompositions with complexity ($O(d_{k}^{3})$) for $k\in\{a,b\}$ that we explicitly aim to avoid. Instead, we analyze the stationarity conditions (˜6, Appx.˜F) and derive a new variant, idealized F-Shampoo (Fig.˜6, Appx.˜F), that is structurally similar to KL-Shampoo. While a straightforward implementation of F-Shampoo performs poorly in practice, the techniques we develop for KL-Shampoo in Sec.˜4 can be adapted to improve its performance (see Fig.˜7, LABEL:{app:extra_exp}).
von Neumann divergence
@@
Another Shampoo variant, often discussed in the literature (Morwani et al., 2025; Vyas et al., 2025a; Eschenhagen et al., 2025), is Shampoo with trace scaling. Vyas et al. (2025a) established that Shampoo with trace scaling is equivalent to running Adafactor (Shazeer & Stern, 2018) in Shampoo’s eigenbasis. In contrast, KL-Shampoo is not equivalent to running Adafactor in its eigenbasis. To clarify this distinction, we make the theoretical connection between Shampoo and Adafactor more explicit: Shampoo with trace scaling is exactly a matrix generalization of Adafactor obtained by minimizing the von Neumann (VN) divergence (Tsuda et al., 2005; Dhillon & Tropp, 2008; Nock et al., 2012) and recovers Adafactor when its Kronecker factors are restricted to be diagonal, as we establish in ˜7 (Appx.˜G). By contrast, KL-Shampoo minimizes the KL divergence instead of the VN divergence. A straightforward implementation of Shampoo with trace scaling—referred to as idealized VN-Shampoo—performs poorly in practice. Howeve, the techniques we develop for KL-Shampoo in Sec.˜4 can be adapted (Fig.˜8, Appx.˜G) to substantially improve its performance, as shown in Fig.˜10 (LABEL:{app:extra_exp}).
@@
A natural question then arises: which “distance” is more suitable? Theoretically, the KL divergence is broadly applicable to arbitrary SPD matrices (Bhatia, 2007; Boumal et al., 2014) and is widely used for covariance matrices (Minh & Murino, 2017). In contrast, the Frobenius norm does not respect the SPD constraint, and the VN divergence is primarily motivated, studied, and applied for unit-trace SPD matrices (Tsuda et al., 2005; Nielsen & Chuang, 2010). Empirically, adopting the KL divergence yields larger improvements than both the Frobenius norm and the VN divergence for designing Shampoo’s schemes (see Fig.˜5) and in other applications (Kulis et al., 2009).
@@ 5 Interpreting and Improving SOAP via KL Minimization
Recall that SOAP (Eq.˜2) applies Shampoo’s scheme to estimate its Kronecker factors and then performs RMSProp updates in the eigenbasis of these factors. Consequently, the interpretation of SOAP’s Kronecker factor estimation is identical to that of Shampoo. RMSProp’s second-moment estimation in the eigenbasis can itself be interpreted as the optimal solution to a separate KL divergence minimization problem, as established in ˜5 (see Appx.˜E for a proof). The KL perspective—distinct from the Frobenius-norm viewpoint (George et al., 2018; Eschenhagen et al., 2025)—provides a new lens for understanding RMSProp’s estimation in the eigenbasis as the estimation of augmented eigenvalues of a covariance matrix under KL divergence. When an outdated eigenbasis is used, RMSProp’s scheme (Step 4a of Fig.˜3) for eigenvalue estimation can be viewed as a correction in an augmented (full-diagonal) space, ${\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$, analogous in spirit to the Frobenius-norm interpretation but derived under the KL framework. This perspective also highlights a close similarity to KL-Shampoo’s estimation scheme: recall that we introduced a comparable correction (Step 3a of Fig.˜3) for KL-Shampoo, but in the original Kronecker-factored diagonal space, ${\bm{Q}}\mathrm{Diag}({\bm{\lambda}}_{a}\otimes{\bm{\lambda}}_{b}){\bm{Q}}^{\top}$.

### v4

@@ 1 Introduction
In this work, we introduce a KL perspective that interprets the estimation schemes of Shampoo and SOAP as solutions to KL-minimization problems for covariance estimation. Our approach naturally extends to tensor-valued settings, where some existing theoretical interpretations may not apply. This perspective reveals a key limitation (illustrated in Fig.˜5): the Kronecker-structured estimators used by Shampoo and SOAP do not adequately solve the corresponding KL-minimization problem. This limitation, in turn, opens new opportunities for improvement. Leveraging this insight, we refine the estimation rules of Shampoo and SOAP and develop practical KL-based schemes—KL-Shampoo and KL-SOAP—that meet or exceed the performance of Shampoo and SOAP for NN (pre-)training while maintaining SOAP-level per-iteration runtime. Notably, KL-Shampoo does not rely on Adam to achieve competitive performance, thereby avoiding Adam’s additional memory overhead (Footnote˜1). In addition, our practical techniques developed for KL-Shampoo (Sec.˜4) can be adapted to strengthen existing Shampoo variants and make the trace-scaling variant competitive without step-size grafting (Fig.˜10, Appx.˜H), while also achieving SOAP-level per-iteration runtime (Fig.˜10, Appx.˜H). Empirically, we show that KL-based methods are competitive for training a range of NNs and remain as flexible as Shampoo and SOAP for tensor-valued weights. Surprisingly, KL-Shampoo consistently outperforms the other methods in our experiments (Figs.˜1 and 5). Overall, our KL-based approach provides a principled way for designing structured methods in NN optimization.
@@ 2 Background
where $\hat{{\bm{g}}}:={\bm{Q}}^{\top}{\bm{g}}=\operatorname{vec}({\bm{Q}}_{a}^{\top}{\bm{G}}{\bm{Q}}_{b})$ is a “projected” gradient vector in eigenbasis $\bm{Q}$} and recall that ${\bm{S}}:={\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$ is SOAP’s preconditioner. Here, we leverage the Kronecker structure and orthogonality of the eigenbasis to move from the left to the right in the last line of Eq.˜2. Note that this EMA weight $\beta_{2}$ is defined as $1-\beta_{2}^{\text{(Adam)}}$, where $\beta_{2}^{\text{(Adam)}}$ is Adam’s (RMSProp’s) $\beta_{2}$. We use this definition rather than Adam’s because we want to further interpret this moving-average scheme through the lens of our KL perspective.
@@ 3 Second Moment Estimation via Kullback–Leibler Minimization
where $\smash{\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]}$ and $\bm{S}$} are considered as Gaussian’s covariance, $\mathrm{det}(\cdot)$ denotes the matrix determinant of its input, and $\kappa\!\geq\!0$ is a damping weight to ensure the positive-definiteness of $\smash{\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]\!+\!\kappa{\bm{I}}}$ if necessary. Mathematically, this KL divergence coincides (up to a factor of $\nicefrac{{1}}{{2}}$) with the log-determinant divergence widely used in matrix optimization (Dhillon & Tropp, 2008; Kulis et al., 2009; Sra, 2016) for SPD matrices rather than Gaussian distributions, which does not require a zero-mean assumption for SPD matrices. This additional zero-mean Gaussian viewpoint provides a probabilistic interpretation of this SPD-aware “distance”, even when the target matrix is not itself a second moment, such as the curvature matrix used in quasi-Newton methods (Fletcher, 1991; Waldrip & Niven, 2016). Moreover, the KL divergence naturally extends to tensor-valued cases, such as a 3D tensor gradient, ${\bm{G}}\in\mathcal{R}^{d_{a}\times d_{b}\times d_{c}}$, by considering a structured preconditioner ${\bm{S}}={\bm{S}}_{a}\otimes{\bm{S}}_{b}\otimes{\bm{S}}_{c}$ to approximate the flattened gradient second moment, $\smash{\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]}$, where matrix ${\bm{S}}_{k}\in\mathcal{R}^{d_{k}\times d_{k}}$ is SPD for $k\in\{a,b,c\}$.
@@
Many works (Morwani et al., 2025; Eschenhagen et al., 2025; An et al., 2025; Xie et al., 2025) primarily focus on matrix-valued weights and interpret Shampoo’s and SOAP’s estimation rules from the Frobenius-norm perspective. However, this norm does not account for the SPD constraint implicitly imposed on Shampoo’s and SOAP’s preconditioners, which ensures that the preconditioned gradient direction is a descent direction (Nesterov et al., 2018). As emphasized in the literature (Pennec et al., 2006; Bhatia, 2007), it is more appropriate to consider a “distance” that respects this constraint. We adopt the KL divergence because it naturally incorporates the SPD constraint, is widely used for covariance estimation (Amari, 2016; Minh & Murino, 2017), and provides a unified framework for reinterpreting and improving Shampoo’s estimation—even in tensor-valued settings where some existing interpretations based on singular value decomposition (SVD) (Van Loan & Pitsianis, 1993) may not apply. In numerical optimization, the KL divergence, known as a merit function (Byrd & Nocedal, 1989), offers a unifying interpretation (Fletcher, 1991; Waldrip & Niven, 2016) and extension (Kanamori & Ohara, 2013a, b) of the low-rank estimation schemes of BFGS and DFP. In contrast, the standard Frobenius norm cannot recover these updates without additional weighting (see Sec. 6.1 of Nocedal & Wright (2006)). In statistical estimation and inference, this KL divergence is also preferred over the Frobenius norm (James & Stein, 1961; Kivinen & Warmuth, 1999; Davis & Dhillon, 2006; Khan & Lin, 2017; Lin et al., 2019; Kunstner et al., 2021).
@@ 3.2 Improving Shampoo’s estimation: Idealized KL-Shampoo
˜2 (proof in Appx.˜B) establishes a closed-form condition (see Eq.˜4) when simultaneously learning both Kronecker factors to minimize the KL problem. Statistically, this condition corresponds to the maximum-likelihood covariance estimator of a zero-mean matrix Gaussian (Dutilleul, 1999) when $\mathbb{E}({\bm{g}}{\bm{g}}^{\top})$ is considered as a finite average $\frac{1}{N}\sum_{i=1}^{N}{\bm{g}}_{i}{\bm{g}}_{i}^{\top}$. This is because $\mathrm{KL}(\frac{1}{N}\sum_{i=1}^{N}{\bm{g}}_{i}{\bm{g}}_{i}^{\top},{\bm{S}})=-\frac{1}{N}\sum_{i=1}^{N}\log\mathcal{N}({\bm{g}}_{i};0,{\bm{S}})+\text{const}$, where ${\bm{g}}_{i}$ is considered as a sample generated from $\mathcal{N}(0,{\bm{S}})$. Thus, satisfying this condition immediately implies that gradient matrix $\bm{G}$} is generated from a zero-mean matrix Gaussian, ${\bm{G}}\sim\mathcal{MN}(\mathbf{0},{\bm{S}}_{a}^{*},{\bm{S}}_{b}^{*})$, with row-wise covariance ${\bm{S}}_{a}^{*}$ and column-wise covariance ${\bm{S}}_{b}^{*}$ obtained by maximum likelihood. Under this condition, Shampoo-style preconditioning naturally induces matrix-Gaussian (row and column) whitening: $\big({\bm{S}}_{a}^{*}\big)^{-1/2}{\bm{G}}\big({\bm{S}}_{b}^{*}\big)^{-1/2}\sim\mathcal{MN}(\mathbf{0},{\bm{I}}_{a},{\bm{I}}_{b})$. As we will discuss in Sec.˜5, the SOAP-like projection onto the optimal eigenbasis leads to covariance diagonalization. Our KL-based approach further extends this to tensor-Gaussian whitening for tensor-valued gradients—without the prohibitive cost typically associated with SVD-based methods. Notably, this kind of whitening arises from minimizing KL divergence, not the Frobenius norm. In machine learning, Lin et al. (2019, 2024) treated this condition as a theoretical example of a multilinear exponential-family (see Sec 5 of Lin et al. (2019)) for Kronecker-based optimization via natural gradient descent on a matrix Gaussian manifold, while more recently, Vyas et al. (2025b) considered a similar condition motivated heuristically by gradient whitening. However, we cannot directly use this condition due to the correlated update between ${\bm{S}}_{a}^{*}$ and ${\bm{S}}_{b}^{*}$. For example, solving ${\bm{S}}_{a}^{*}$ requires knowing ${\bm{S}}_{b}^{*}$ in Eq.˜4 or vice versa. In practice, this condition is unachievable because the expectations in Eq.˜4 must be approximated. Thus, we consider an estimated ${\bm{S}}_{k}$ to approximate ${\bm{S}}_{k}^{*}$ for $k\in\{a,b\}$ and propose an exponential moving average (EMA) scheme:
@@ 3.3 Comparison with Kronecker Schemes Using Alternative Divergences
Morwani et al. (2025) consider a two-sided Shampoo variant based on the Frobenius norm and derive the optimal solution via rank-1 singular value decomposition (SVD) of the second moment $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ (Van Loan & Pitsianis, 1993). However, this solution is often unattainable in practice and is computationally expensive for two reasons: (1) the expectation $\mathbb{E}[{\bm{g}}{\bm{g}}^{\top}]$ must be approximated; and (2) performing the SVD is costly—yielding complexity ($O(d_{a}^{2}d_{b}^{2})$) in general even for rank-1 SVD—which is higher than the eigen decompositions with complexity ($O(d_{k}^{3})$) for $k\in\{a,b\}$ that we aim to avoid. Instead, we analyze the stationarity conditions (˜6, Appx.˜F) and derive a new variant, idealized F-Shampoo (Fig.˜6, Appx.˜F), that is structurally similar to KL-Shampoo. While a straightforward implementation of F-Shampoo performs poorly in practice, the techniques (Sec.˜4) we develop for KL-Shampoo can be adapted to improve its performance (Fig.˜7, LABEL:{app:extra_exp}).
Von Neumann divergence (VN-Shampoo)
@@
Another variant, often discussed in the literature (Morwani et al., 2025; Vyas et al., 2025a; Eschenhagen et al., 2025), is Shampoo with trace scaling. Vyas et al. (2025a) established that Shampoo with trace scaling is equivalent to running Adafactor (Shazeer & Stern, 2018) in Shampoo’s eigenbasis. In contrast, KL-Shampoo is not equivalent to running Adafactor in its eigenbasis. To clarify this distinction, we make the theoretical connection between Shampoo and Adafactor more explicit: Shampoo with trace scaling is exactly a matrix generalization of Adafactor obtained by minimizing the von Neumann (VN) divergence (Tsuda et al., 2005; Dhillon & Tropp, 2008; Nock et al., 2012) and recovers Adafactor when its Kronecker factors are restricted to be diagonal, as we establish in ˜7 (Appx.˜G). By contrast, KL-Shampoo minimizes the KL divergence instead of the VN divergence. A straightforward implementation of Shampoo with trace scaling—referred to as idealized VN-Shampoo—performs poorly in practice. Howeve, the techniques we develop for KL-Shampoo in Sec.˜4 can be adapted (Fig.˜8, Appx.˜G) to substantially improve its performance so that it matches Shampoo with step-size grafting and outperforms SOAP, as shown in Figs.˜10 and 10 (LABEL:{app:extra_exp}).
@@
A natural question then arises: which divergence is more suitable? Theoretically, the KL divergence is broadly applicable to arbitrary SPD matrices (Bhatia, 2007; Boumal et al., 2014) and is widely used for covariance matrices (Minh & Murino, 2017). In contrast, the Frobenius norm does not respect the SPD constraint, and the VN divergence is primarily motivated, studied, and applied for unit-trace SPD matrices (Tsuda et al., 2005; Nielsen & Chuang, 2010). Empirically, adopting the KL divergence yields larger improvements than both the Frobenius norm and the VN divergence for designing Shampoo’s schemes (see Fig.˜5) and in other applications (Kulis et al., 2009).
@@ 5 Interpreting and Improving SOAP via KL Minimization
Recall that SOAP (Eq.˜2) applies Shampoo’s scheme to estimate its Kronecker factors and then performs RMSProp updates in the eigenbasis of these factors. Consequently, the interpretation of SOAP’s Kronecker factor estimation is identical to that of Shampoo. RMSProp’s second-moment estimation in the eigenbasis can itself be interpreted as the optimal solution to a separate KL divergence minimization problem, as established in ˜5 (see Appx.˜E for a proof). The KL perspective—distinct from the Frobenius-norm viewpoint (George et al., 2018; Eschenhagen et al., 2025)—provides a new lens for understanding RMSProp’s estimation in the eigenbasis as the estimation of augmented eigenvalues of a covariance matrix under KL divergence. When an outdated eigenbasis is used, RMSProp’s scheme (Step 4a of Fig.˜3) for eigenvalue estimation can be viewed as a correction in an augmented (full-diagonal) eigen space, ${\bm{Q}}\mathrm{Diag}({\bm{d}}){\bm{Q}}^{\top}$, analogous in spirit to the Frobenius-norm interpretation but derived under the KL framework. This perspective also highlights a close similarity to KL-Shampoo’s estimation scheme: recall that we introduced a comparable correction (Step 3a of Fig.˜3) for KL-Shampoo, but in the original Kronecker-factored diagonal eigen space, ${\bm{Q}}\mathrm{Diag}({\bm{\lambda}}_{a}\otimes{\bm{\lambda}}_{b}){\bm{Q}}^{\top}$. Notably, when using KL-Shampoo’s eigenbasis ${\bm{Q}}^{*}={\bm{Q}}_{a}^{*}\otimes{\bm{Q}}_{b}^{*}$ obtained from the optimal condition in Eq.˜4, we can see the (Gaussian) covariance of gradient vector $\bm{g}$} in the eigenbasis is Kronecker-diagonalized rather than fully diagonalized: $\mathrm{vec}(({\bm{Q}}_{a}^{*})^{\top}{\bm{G}}{\bm{Q}}_{b}^{*})=({\bm{Q}}^{*})^{\top}{\bm{g}}\sim\mathcal{N}(0,\mathrm{Diag}({\bm{\lambda}}_{a}^{*})\otimes\mathrm{Diag}({\bm{\lambda}}_{b}^{*}))$.